Find the measure of an angle whose supplement measures 43 degrees more than twice its complement
Answer by mananth(16949)

let the angle be x deg
supplement will be 180-x
complement will be 90-x
..
supplement = 2(complement )+43
..
180-x=2(90-x)+43
180-x=180-2x+43
2x-x=43
x=43 deg